# Pushing the Limits of Sparsity: A Bag of Tricks for Extreme Pruning

### v3

@@ 4.1 Complexity Analysis
While DCTpS achieves higher accuracy than standard DST methods at extreme sparsity levels, it comes with substantial computational overhead that limits its practical deployment benefits. DCTpS uses a dense DCT matrix, which theoretically could have a complexity of . We found that its theoretical inference FLOPs are about 32X higher than the standard DST at 99.95% and 150X higher at 99.99%. This is because the dense DCT computation dominates at extreme sparsities, becoming an increasingly significant bottleneck as sparsity increases. This additional computational overhead directly explains the inference throughput and latency differences observed in our experiments.
@@ 4.2 Ablation Studies
Improving gradient flow during early training has been shown to be critical for successfully training sparse networks. Evci et al. [5] demonstrated that methods with better gradient flow, particularly during initialization and early training, consistently achieved superior performance. Building on this insight, we investigated the effect of DyReLU phasing on gradient flow in addition to measuring accuracy.
@@
To analyze this, we computed the sum of the gradient norm of all layers, and compared the gradient flow during training between models using DyReLU phasing against those with standard ReLU. As shown in the top row of Figure 5, models using only standard ReLU have zero gradient norms at 99.99% sparsity and this is reflected in their collapsed accuracy. On the other hand, the networks with DyReLU phasing receive high gradient flow in the beginning where , during phasing where , and maintains healthy gradient flow when . Results suggest that the network retains some of the gradient-boosting benefits initially provided by DyReLU, We conjecture that this is because DyReLU amplifies important features in the beginning and helps the network settle into a configuration that preserves gradients even after switching to standard ReLU. Gradient flow especially during early stages is a strong indicator of the DST’s success, and helps explain the performance gap between the methods. This supports our theoretical findings in the theorem, confirming that DyReLU phasing indeed provides greater gradient flow than the standard ReLU.
@@
Similarly, we isolate the effect of weight sharing by adding only this mechanism to the RigL pipeline, keeping all other settings the same. Once again, gradient flow has been restored and we see a dramatic increase in gradient flow, as shown in the bottom row of Figure 5. This supports our hypothesis that creating additional gradient paths through weight sharing can help maintain stable training at extreme sparsities.
Effect of Cyclic Sparsity. Cyclic sparsity provides a flexible framework that allows networks to explore a wider parameter space. We perform an ablation study on the effect of cyclic sparsities using ResNet-34 and CIFAR-100 with a target sparsity of 99.95%. As shown in Figure 6, different cyclic schedules - whether starting from high sparsity and decreasing, or from low sparsity and increasing, consistently outperform the static sparsity (red) across all configurations. While each schedule can be tuned differently (e.g., altitude, length of cycle, etc.), we consistently observe improved parameter exploration and generalization. We conclude that having some form of cyclic sparsity may improve the performance over static sparsity. However, the large number of possible configurations makes systematic experimentation challenging. Future work will focus on tuning these schedules to identify optimal schedules for different datasets and architectures.
@@ 5 Conclusion
In this paper, we introduced EAST, a new DST approach focusing on extreme sparsities that incorporates DyReLU phasing, weight sharing, and cyclic sparsity scheduling. Through extensive empirical evaluation on CIFAR-10, CIFAR-100 and ImageNet using ResNet-34 and ResNet-50, we demonstrated EAST’s ability to achieve competitive performance at extreme sparsities, surpassing existing methods by a large margin. Each component of EAST individually contributes to maintaining robust gradient flow and optimizing parameter exploration at extreme sparsity, while their combined effect can further enhance performance stability. Our results highlight the potential for practical applicability in resource-constrained environments, establishing it as a scalable and effective method for highly compressed networks. This work opens up future research to optimize models at extreme sparsities, an area often overlooked in previous studies.
Acknowledgments. Our work was supported by the UKRI EPSRC funding council (EP/V042270/1), a University of Aberdeen PhD studentship, the HPC facility ”Maxwell”, and the Interdisciplinary Institute.
Supplementary Material
@@ A.2 Dynamic ReLU
A.3 Phasing out DyReLU
As training progresses, we phase out all DyReLU and gradually replace them with the standard ReLU. We introduce a hyperparameter that regulates the contribution of DyReLU during training. It makes full contribution when it is 1, and gradually decreases to 0 as training progresses, eventually reaching 0 before the first learning rate decay, at which point DyReLU has been completely replaced by ReLU. We find that this helps the network settle into a more stable configuration, as DyReLU is no longer needed once the network has explored the parameter space adequately. This process removes all DyReLU-introduced parameters.
By the end of phasing , we have:
@@ B. Implementation Details
| ResNet-34, | EAST | CIFAR-10/100 | 128 | 250 | 0.1 | 10x[75,150] | SGD(0.9) | 1.0e-4 | ERK | 2 | 50 | 3 | 350 | 4000 |

### v4

@@ 4.1 Complexity Analysis
While DCTpS achieves higher accuracy than standard DST methods at extreme sparsity levels, it incurs substantial computational overhead, limiting its practical deployment benefits. DCTpS uses a dense DCT matrix, which theoretically could have a complexity of . We found that its theoretical inference FLOPs are about 32X higher than the standard DST at 99.95% and 150X higher at 99.99%. This is because the dense DCT computation dominates at extreme sparsities, becoming an increasingly significant bottleneck as sparsity increases. This additional computational overhead directly explains the inference throughput and latency differences observed in our experiments.
@@ 4.2 Ablation Studies
Improving gradient flow during early training has been shown to be critical for successfully training sparse networks. Evci et al. (2022) demonstrated that methods with better gradient flow, particularly during initialization and early training, consistently achieved superior performance. Building on this insight, we investigated the effect of DyReLU phasing on gradient flow in addition to measuring accuracy.
@@
To analyze this, we computed the sum of the gradient norm of all layers, and compared the gradient flow during training between models using DyReLU phasing against those with standard ReLU. As shown in the top row of Figure 5, models using only standard ReLU have zero gradient norms at 99.99% sparsity and this is reflected in their collapsed accuracy. On the other hand, the networks with DyReLU phasing receive high gradient flow in the beginning where , during phasing where , and maintains healthy gradient flow when . Results suggest that the network retains some of the gradient-boosting benefits initially provided by DyReLU, We conjecture that this is because DyReLU amplifies important features in the beginning and helps the network settle into a configuration that preserves gradients even after switching to standard ReLU. Gradient flow especially during early stages is a strong indicator of the DST’s success, and helps explain the performance gap between the methods.
@@
Similarly, we isolate the effect of weight sharing by adding only this mechanism to the RigL pipeline, keeping all other settings the same. Once again, gradient flow has been restored, and we see a dramatic increase in gradient flow, as shown in the bottom row of Figure 5. This supports our hypothesis that creating additional gradient paths through weight sharing can help maintain stable training at extreme sparsities.
Effect of Cyclic Sparsity. We evaluate the impact of dynamically varying sparsity levels throughout training compared to maintaining static sparsity. Using ResNet-34 and ResNet-50 on CIFAR-100 at 99.95% target sparsity, we test different cyclic schedules against static sparsity baselines. As shown in Figure 6, cyclic sparsity consistently outperforms static sparsity across both architectures to varying degrees. The cyclic schedule allows the sparsity to temporarily relax before gradually returning to the target extreme level. This periodic increase in connectivity encourages parameter exploration during the denser phases while maintaining the desired final sparsity level. While ResNet-34 shows incremental improvement with cyclic scheduling, ResNet-50 demonstrates more substantial gains, suggesting that deeper networks may benefit more from this cyclic approach.
We provide more test results using different cyclic hyperparameters in Appendix C. Based on systematic evaluation across multiple configurations (Supplementary Tables C.2-C.5), we observe several trends. 10 density multipliers generally outperform 3, but 30 is not necessarily better than 10, suggesting there may be an elbow point where further increases no longer improves performance. Having two cycles generally performs better than having one, with longer subsequent cycle providing additional gains. While optimal configurations may vary by architecture and dataset, we recommend starting with a 10 multiplier and a length multiplier of 2 if two cycles are used when tuning cyclic sparsity hyperparameters.
@@ 5 Conclusion
In this paper, we introduced EAST, a new DST approach that focuses on extreme sparsity and incorporates DyReLU phasing, weight sharing, and cyclic sparsity scheduling. Through extensive empirical evaluation on CIFAR-10, CIFAR-10, and ImageNet using ResNet-34 and ResNet-50, we demonstrated that EAST achieves competitive performance at extreme sparsitys, surpassing existing methods by a large margin. Each component of EAST individually contributes to maintaining robust gradient flow and optimizing parameter exploration at extreme sparsity, while their combined effect can further enhance performance stability. Our results highlight the potential for practical applicability in resource-constrained environments, establishing it as a scalable and effective method for highly compressed networks. This work opens up future research to optimize models at extreme sparsities, an area often overlooked in previous studies.
Acknowledgments: Our work was supported by the UKRI EPSRC funding council (EP/V042270/1), a University of Aberdeen PhD studentship, the HPC facility ”Maxwell”, and the Interdisciplinary Institute.
Appendix A Supplementary Material
@@ B. Implementation Details
| ResNet-34, | EAST | CIFAR-10/100 | 128 | 250 | 0.1 | 10x[75,150] | SGD(0.9) | 1.0e-4 | ERK | 2 | 50 | 10 | 350 | 4000 |
@@
C. Cyclic Sparsity Analysis
We tested two cyclic patterns - cosine schedules that follow smooth cosine curves and triangular schedules that transition linearly between top and bottom sparsity levels. The key hyperparameters are: the density multiplier (how many times the peak density exceeds the target sparsity - e.g., with a target density of 0.0005, a 3 multiplier peaks at 0.0015 while 10 peaks at 0.005), the number of cycles, and the cycle’s length multiplier (how many times longer each subsequent cycle is compared to the previous one). As shown in Tables C.2-C.5, cyclic sparsity can be sensitive to hyperparameter choices and sometimes yields inconsistent results, particularly with ResNet-34 where certain configurations (e.g., cosine patterns with low multipliers) may decrease performance compared to static sparsity. This suggests that cyclic schedules may require careful tuning to be effective. However, several consistent trends emerge across our experiments: (1) a higher density multiplier of 10 generally outperform a lower one of 3, (2) having two cycles tend to be more effective than just one, and (3) longer subsequent cycles often improve performance vs two cycles of equal length. While these patterns suggest that dynamic sparsity variation has merit, the high sensitivity to hyperparameters and occasional performance degradation indicate that cyclic sparsity should be considered a more experimental component of our framework, requiring further research to develop principled guidelines for hyperparameter selection.
| Method | Pattern | Multiplier | Cycles | Length Mult. | Test Accuracy (%) | Change (%) |
| Static | – | – | – | – | 81.470.47 | – |
| Cyclic | Cosine | 3 | 1 | - | 83.570.26 | +2.1 |
| Cyclic | Cosine | 3 | 2 | 1 | 83.980.14 | +2.5 |
| Cyclic | Cosine | 3 | 2 | 2 | 84.080.31 | +2.6 |
| Cyclic | Cosine | 10 | 1 | - | 84.270.13 | +2.8 |
| Cyclic | Cosine | 10 | 2 | 1 | 85.210.19 | +3.7 |
| Cyclic | Cosine | 10 | 2 | 2 | 85.340.18 | +3.9 |
| Cyclic | Cosine | 30 | 2 | 1 | 84.890.24 | +3.4 |
| Cyclic | Cosine | 30 | 2 | 2 | 84.940.39 | +3.5 |
| Cyclic | Triangular | 3 | 1 | - | 82.980.09 | +1.5 |
| Cyclic | Triangular | 3 | 2 | 1 | 84.160.06 | +2.7 |
| Cyclic | Triangular | 3 | 2 | 2 | 84.550.34 | +3.1 |
| Cyclic | Triangular | 10 | 1 | - | 84.320.15 | +2.8 |
| Cyclic | Triangular | 10 | 2 | 1 | 86.100.24 | +4.6 |
| Cyclic | Triangular | 10 | 2 | 2 | 85.460.04 | +4.0 |
| Cyclic | Triangular | 30 | 2 | 1 | 85.790.25 | +4.3 |
| Cyclic | Triangular | 30 | 2 | 2 | 85.510.14 | +4.0 |
| Method | Pattern | Multiplier | Cycles | Length Mult. | Test Accuracy (%) | Change (%) |
| Static | – | – | – | – | 38.440.71 | – |
| Cyclic | Cosine | 3 | 1 | - | 34.370.23 | -4.1 |
| Cyclic | Cosine | 3 | 2 | 1 | 37.170.81 | -1.3 |
| Cyclic | Cosine | 3 | 2 | 2 | 38.360.48 | -0.1 |
| Cyclic | Cosine | 10 | 1 | - | 38.680.35 | +0.2 |
| Cyclic | Cosine | 10 | 2 | 1 | 41.350.54 | +2.9 |
| Cyclic | Cosine | 10 | 2 | 2 | 44.831.23 | +6.4 |
| Cyclic | Cosine | 30 | 2 | 1 | 42.451.04 | +4.0 |
| Cyclic | Cosine | 30 | 2 | 2 | 43.830.23 | +4.4 |
| Cyclic | Triangular | 3 | 1 | - | 34.631.52 | -3.8 |
| Cyclic | Triangular | 3 | 2 | 1 | 36.491.87 | -2.0 |
| Cyclic | Triangular | 3 | 2 | 2 | 38.801.10 | +0.4 |
| Cyclic | Triangular | 10 | 1 | - | 41.250.51 | +2.8 |
| Cyclic | Triangular | 10 | 2 | 1 | 43.970.74 | +5.5 |
| Cyclic | Triangular | 10 | 2 | 2 | 45.280.68 | +6.8 |
| Cyclic | Triangular | 30 | 2 | 1 | 44.150.24 | +5.7 |
| Cyclic | Triangular | 30 | 2 | 2 | 44.890.56 | +6.5 |
| Method | Pattern | Multiplier | Cycles | Length Mult. | Test Accuracy (%) | Change (%) |
| Static | – | – | – | – | 30.1510.85 | – |
| Cyclic | Cosine | 3 | 1 | - | 75.172.87 | +45.0 |
| Cyclic | Cosine | 3 | 2 | 1 | 75.673.19 | +45.5 |
| Cyclic | Cosine | 3 | 2 | 2 | 76.443.67 | +46.3 |
| Cyclic | Cosine | 10 | 1 | - | 78.754.06 | +48.6 |
| Cyclic | Cosine | 10 | 2 | 1 | 78.304.70 | +48.2 |
| Cyclic | Cosine | 10 | 2 | 2 | 77.374.87 | +47.2 |
| Cyclic | Cosine | 30 | 2 | 1 | 79.252.70 | +49.1 |
| Cyclic | Cosine | 30 | 2 | 2 | 79.681.79 | +49.5 |
| Cyclic | Triangular | 3 | 1 | - | 76.083.14 | +45.9 |
| Cyclic | Triangular | 3 | 2 | 1 | 75.064.40 | +44.9 |
| Cyclic | Triangular | 3 | 2 | 2 | 78.294.51 | +48.1 |
| Cyclic | Triangular | 10 | 1 | - | 78.555.29 | +48.4 |
| Cyclic | Triangular | 10 | 2 | 1 | 78.275.32 | +48.1 |
| Cyclic | Triangular | 10 | 2 | 2 | 79.234.93 | +49.1 |
| Cyclic | Triangular | 30 | 2 | 1 | 77.653.24 | +47.5 |
| Cyclic | Triangular | 30 | 2 | 2 | 79.014.32 | +48.9 |
| Method | Pattern | Multiplier | Cycles | Length Mult. | Test Accuracy (%) | Change (%) |
| Static | – | – | – | – | 7.024.63 | – |
| Cyclic | Cosine | 3 | 1 | - | 33.711.57 | +26.7 |
| Cyclic | Cosine | 3 | 2 | 1 | 34.323.91 | +27.3 |
| Cyclic | Cosine | 3 | 2 | 2 | 35.343.95 | +28.3 |
| Cyclic | Cosine | 10 | 1 | - | 39.193.73 | +32.2 |
| Cyclic | Cosine | 10 | 2 | 1 | 39.612.83 | +32.6 |
| Cyclic | Cosine | 10 | 2 | 2 | 42.781.47 | +35.8 |
| Cyclic | Cosine | 30 | 3 | 1 | 36.682.95 | +29.7 |
| Cyclic | Cosine | 30 | 3 | 2 | 41.791.54 | +34.8 |
| Cyclic | Triangular | 3 | 1 | - | 36.192.20 | +29.2 |
| Cyclic | Triangular | 3 | 2 | 1 | 36.583.06 | +29.6 |
| Cyclic | Triangular | 3 | 2 | 2 | 37.613.21 | +30.6 |
| Cyclic | Triangular | 10 | 1 | - | 42.073.83 | +35.0 |
| Cyclic | Triangular | 10 | 2 | 1 | 43.551.57 | +36.5 |
| Cyclic | Triangular | 10 | 2 | 2 | 42.105.70 | +35.1 |
| Cyclic | Triangular | 30 | 2 | 1 | 41.952.57 | +34.9 |
| Cyclic | Triangular | 30 | 2 | 2 | 42.114.80 | +35.1 |
D. Training Cost Analysis
| Method | Train Time (hours) | Param Size (MB) | Est. Total (MB) |
| RigL | <1 | 81.36 | 100.44 |
| DCTpS | 2 | 162.66 | 197.36 |
| EAST w/ DyReLU | 1 | 81.36 | 107.81 |
| EAST w/ WS | <1 | 42.83 | 58.78 |
| EAST w/ Cyclic | <1 | 81.36 | 100.44 |
| EAST Combined | 1 | 54.96 | 78.78 |
EAST adds minimal training overhead compared to standard DST (RigL), while DCTpS requires 2× training time due to dense offset matrix computations.
DyReLU phasing: Introduces additional parameters during early training (7 MB activation overhead), which are removed after phasing completes.
Weight sharing: Directly reduces stored parameters by eliminating redundant weights across residual blocks (81→43 MB), providing immediate memory savings.
Cyclic sparsity: Memory footprint identical to baseline RigL, as it only modifies the mask update schedule without additional parameters or activations.